In a right triangle, the length of one leg is two less than of the other leg. If the length of the hypotenuse is equal to the square root of 34, find the lengths of the two legs?
Answer by edjones(7569)

a=b+2, c=sqrt(34)
a^2+b^2=c^2
(b+2)^2+b^2=34
b^2+4b+4+b^2=34
2b^2+4b-30=0
b^2+2b-15=0
(b+5)(b-3)=0
b=3, a=5 or vice versa.
.
Ed